This one has kicked my skunk over and over! I know the answer is 4 & 20, but I cannot see how to arrive at that answer.
4y+20 under the radical sign - y-4 under the radical sign = 6
I hope I explained that well enough. I tried to cut and paste the problem out of my Microsoft Word file but it did not work.

Answer by stanbon(75887)

4y+20 under the radical sign - y-4 under the radical sign = 6
---------------------------------
Rearrange:
sqrt(4y+20) = sqrt(y-4)+6
Square both sides to get:
(4y+20) = (y-4)+12sqrt(y-4)+36
12sqrt(y-4)= 4y+20-(y-4)-36
12sqrt(y-4) = 3y-12
4sqrt(y-4) = y-4
Square both sides to get:
16(y-4) = y^2-8y+16
y^2-24y+80 =0
(y-4)(y-20)=0
y = 4 or y = 20
===================
Check these answers in the original equation.
=====================
Cheers,
Stan H.